# Reasoning by Exploration: A Unified Approach to Retrieval and Generation over Graphs

### v1

@@ 4.2. Stage I: Supervised Fine-Tuning for Step-wise Exploration
In the first stage, we aim to equip the LLM with the basic ability to perform step-wise graph exploration. Specifically, we require the gold actions for each step, which include both the step-wise answers and the new exploration paths as defined in Section 4.1. Previous works (Zhang et al., 2022b; Sun et al., 2019) typically use the shortest path from the query entities to the answers to supervise their models. However, we argue that relying solely on the shortest path is overly restrictive. In real knowledge graphs, there may exist multiple valid reasoning paths connecting the same query entity to the correct answer; constraining training to only the shortest one discourages exploration and limits generalization. Moreover, the model does not have a global view of the entire graph at each step. When several neighboring entities share the same relation with the current frontier entity, selecting only a single target node can be ambiguous, especially when entity names are represented as IDs.
@@
In the supervised fine-tuning (SFT) stage, each training instance is represented as a pair $(s_{d},a_{d}^{*})$ in $\mathcal{D}_{\text{SFT}}$, where $s_{d}$ denotes the state at step $d$ (including the question, current reasoning paths, and local neighborhood), and $a_{d}^{*}$ is the corresponding gold action constructed in Algorithm 1. At each step, the model takes $s_{d}$ as input and predicts action $a_{d}=(\mathcal{A}_{d},\mathcal{P}_{d})$, which includes the step-wise answers and the new exploration paths. The training objective is to maximize the likelihood of the gold action $a_{d}^{*}$ given its state, leading to the loss:
@@
where $\theta$ denotes the model parameters. This loss encourages the model to imitate the expert exploration behavior, deciding which nodes to expand and which entities to output as answers at each step, thereby acquiring the fundamental ability for structured graph exploration before reinforcement-learning refinement in Stage II.
@@ 4.3. Stage II: Reinforcement Learning for Exploration Optimization
Although the supervised fine-tuning (SFT) stage enables the model to imitate gold trajectories, it may cause the LLM to memorize fixed exploration patterns within a single dataset, limiting its ability to generalize to unseen graphs or question types. To address this, we introduce a reinforcement learning (RL) stage that allows the model to interact with the knowledge graph and refine its exploration strategy through feedback rather than imitation.
During this stage, we not only encourage the LLM to generate correct step-wise answers and exploration paths, but also aim to promote broader and more diverse exploration behaviors, enabling the model to discover additional valid reasoning trajectories and answers beyond those seen in the gold data. To enable such flexible and interpretable optimization, we adopt a rule-based reinforcement learning framework, which provides explicit feedback signals without relying on a learned value model. This choice is inspired by recent advances in reasoning LLMs, such as DeepSeek-R1 (Guo et al., 2025b), which have shown that rule-based rewards can effectively guide reasoning and exploration, while simplifying training compared to the original RLHF framework (Ouyang et al., 2022).
@@
A non-zero reward $\mathcal{R}_{\text{format}}$ is given only when the model successfully follows the required format. This encourages the LLM to produce well-structured outputs, facilitating reliable parsing and consistent downstream evaluation during training.
@@
This combined reward jointly encourages the model to produce well-structured outputs, generate correct answers, and discover diverse and valid reasoning paths during reinforcement learning.
We optimize RoE using Group Relative Policy Optimization (GRPO) (Guo et al., 2025b), a variant of Proximal Policy Optimization (PPO) (Schulman et al., 2017). GRPO eliminates the need for an additional value function approximation by using the average reward of multiple sampled outputs as a relative baseline. For each input question $x$, GRPO samples a group of outputs $\{y_{1},y_{2},\ldots,y_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and optimizes the new policy $\pi_{\theta}$ by maximizing the following objective:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Big(r_{i,t}(\theta)\,\hat{A}_{i,t},$
$\quad\mathrm{clip}\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,\hat{A}_{i,t}\Big)-\gamma D_{\text{KL}}\!\big[\pi_{\theta}(\cdot|x)\,\|\,\pi_{\text{ref}}(\cdot|x)\big]\Bigg]$
where
$r_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}\mid x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid x,y_{i,<t})},\qquad\hat{A}_{i,t}=\frac{R_{i}-\mathrm{mean}(\{R_{j}\}_{j=1}^{G})}{\mathrm{std}(\{R_{j}\}_{j=1}^{G})},$
and $\gamma$ is a hyperparameter balancing task-specific rewards and KL-divergence regularization. This formulation encourages the model to prefer outputs with higher relative rewards within each group, stabilizing training while improving sample efficiency. In our case, the rewards $R_{i}$ correspond to the total rule-based rewards $\mathcal{R}_{\text{total}}$.
@@ 5.1. Experimental Settings
Datasets. To evaluate the effectiveness of RoE, we use two widely adopted datasets: WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018), both constructed upon the Freebase knowledge graph (Bollacker et al., 2008). To further assess the generalization ability of GraphRAG methods across different knowledge graphs, we additionally use MetaQA (Zhang et al., 2018), which is built on the WikiMovies KG (Miller et al., 2016), and PathQuestion (Zhou et al., 2018), which uses a Freebase subset with modified relation names. For cross-graph generalization evaluation, we randomly sample 1,000 multi-hop questions from each of these two datasets.
Evaluation Metrics. We follow previous works (He et al., 2021; Guo et al., 2025a; Li et al., 2024) and adopt two evaluation metrics: Hit and F1. Specifically, the Hit measures the proportion of questions for which at least one correct predicted answer, while F1 provides a balanced evaluation of both precision and recall.
Baselines. We compare RoE with a diverse set of baselines, which can be grouped into four categories: (1) LLM-Only: methods that rely solely on the internal knowledge of LLMs to answer questions without using external knowledge graphs. We include Flan-T5-XL (Chung et al., 2024), Alpaca-7B (Taori et al., 2023), Llama-2-7B-Chat (Touvron et al., 2023), Llama-3.1-8B-Instruct (Touvron et al., 2023), and GPT-4.1 (Achiam et al., 2023). (2) GNN-Only: models that use GNNs for both retrieval and reasoning. We include GraftNet (Sun et al., 2018), SR+NSM (Zhang et al., 2022b), and UniKGQA (Jiang et al., 2022). (3) LLMs + GNNs: hybrid approaches that combine LLMs and GNNs for retrieval and generation. For instance, G-Retriever (He et al., 2024) leverages GNNs to enhance LLM reasoning, while GNN-RAG (Mavromatis and Karypis, 2024) employs GNNs for retrieval and LLMs for generation. (4) LLMs + KG: methods that employ LLMs for both retrieval and generation over knowledge graphs. We compare with ToG (Sun et al., 2023), RoG (Luo et al., 2023), KD-CoT (Wang et al., 2023a), and SubGraphRAG (Li et al., 2024).
@@
To ensure a fair comparison, we train each retriever and generator on a single dataset, rather than jointly training across multiple datasets as done in RoG and GNN-RAG. Moreover, to eliminate the influence of backbone model, we use the same Llama-3.1-8B-Instruct model (Touvron et al., 2023) for most LLM-based baselines, including RoG, KD-CoT, SubGraphRAG, G-Retriever, and GNN-RAG.
@@ 5.2. Overall Performance Comparison
  • The LLM-only models, which lack access to external graph information, generally perform worse than methods that incorporate graphs. However, stronger models, such as GPT-4.1, achieve better results, suggesting that large LLMs already possess internal knowledge relevant to the benchmarks. This observation justifies our choice that using the same LLM backbone for baselines to ensure a fair comparison.
@@ 5.4. Ablation Studies
Figure 4. Performance of RoE and its variants pretrained on the WebQSP dataset. Marker shapes denote different model variants, while filled and hollow markers represent Hit and metrics, respectively.

### v2

@@ 4.2. Stage I: Supervised Fine-Tuning for Step-wise Exploration
In the first stage, we aim to equip the LLM with the basic ability to perform step-wise graph exploration. Specifically, we require the gold actions for each step, which include both the step-wise answers and the new exploration paths as defined in Section 4.1. Previous works (Zhang et al., 2022a; Sun et al., 2019) typically use the shortest path from the query entities to the answers to supervise their models. However, we argue that relying solely on the shortest path is overly restrictive. In real knowledge graphs, there may exist multiple valid reasoning paths connecting the same query entity to the correct answer; constraining training to only the shortest one discourages exploration and limits generalization. Moreover, the model does not have a global view of the entire graph at each step. When several neighboring entities share the same relation with the current frontier entity, selecting only a single target node can be ambiguous, especially when entity names are represented as IDs.
@@
In the supervised fine-tuning (SFT) stage, each training instance is represented as a pair $(s_{d},a_{d}^{*})$ in $\mathcal{D}_{\text{SFT}}$, where $s_{d}$ denotes the state at step $d$ (including the question, current reasoning paths, and local neighborhood), and $a_{d}^{*}$ is the corresponding gold action constructed in Algorithm 1. At each step, the model takes $s_{d}$ as input and predicts action $a_{d}=(\mathcal{A}_{d},\mathcal{P}_{d})$, which includes the step-wise answers and the new exploration paths. The training objective is to maximize the likelihood of the gold action $a_{d}^{*}$ given its state, leading to loss:
@@
where $\theta$ denotes the model parameters. This loss encourages the model to imitate the expert exploration behavior, thereby acquiring the fundamental ability for structured graph exploration before reinforcement-learning refinement in Stage II.
@@ 4.3. Stage II: Reinforcement Learning for Exploration Optimization
Although the SFT stage enables the model to imitate gold trajectories, it may cause the LLM to memorize fixed exploration patterns within a single dataset, limiting its ability to generalize to unseen graphs or question types. To address this, we introduce a reinforcement learning (RL) stage that allows the model to interact with the knowledge graph and refine its exploration strategy through feedback rather than imitation. During this stage, we not only encourage the LLM to generate correct step-wise answers and exploration paths, but also aim to promote broader and more diverse exploration behaviors, enabling the model to discover additional valid reasoning trajectories and answers beyond those seen in the gold data. To enable such flexible and interpretable optimization, we adopt a rule-based reinforcement learning framework, which provides explicit feedback signals without relying on a learned value model. This choice is inspired by recent advances in reasoning LLMs, such as DeepSeek-R1 (Guo et al., 2025a), which have shown that rule-based rewards can effectively guide reasoning and exploration, while simplifying training compared to the original RLHF framework (Ouyang et al., 2022).
@@
A non-zero reward $\mathcal{R}_{\text{format}}$ is given only when the model successfully follows the required format.
@@
This combined reward jointly encourages the model to produce well-structured outputs, generate correct answers, and discover diverse and valid reasoning paths during reinforcement learning. We optimize RoE using Group Relative Policy Optimization (GRPO) (Guo et al., 2025a), a variant of Proximal Policy Optimization (PPO) (Schulman et al., 2017).
@@ 5.1. Experimental Settings
Datasets. To evaluate the effectiveness of RoE, we use two widely adopted datasets: WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018), both constructed upon the Freebase knowledge graph (Bollacker et al., 2008). To further assess the generalization ability, we additionally use MetaQA (Zhang et al., 2018), which is built on the WikiMovies KG (Miller et al., 2016), and PathQuestion (Zhou et al., 2018), which uses a Freebase subset with modified relation names.
Evaluation Metrics. We follow previous works (He et al., 2021; Guo et al., 2025b; Li et al., 2024) and adopt two evaluation metrics: Hit and F1. Specifically, the Hit measures the proportion of questions for which at least one correct predicted answer, while F1 provides a balanced evaluation of both precision and recall.
Baselines. We compare RoE with a diverse set of baselines, which can be grouped into four categories: (1) LLM-Only: methods that rely solely on the internal knowledge of LLMs to answer questions without using external knowledge graphs. We include Flan-T5-XL (Chung et al., 2024), Alpaca-7B (Taori et al., 2023), Llama-2-7B-Chat (Touvron et al., 2023), Llama-3.1-8B-Instruct (Touvron et al., 2023), and GPT-4.1 (Achiam et al., 2023). (2) GNN-Only: models that use GNNs for both retrieval and reasoning. We include GraftNet (Sun et al., 2018), SR+NSM (Zhang et al., 2022a), and UniKGQA (Jiang et al., 2022). (3) LLMs + GNNs: hybrid approaches that combine LLMs and GNNs for retrieval and generation. For instance, G-Retriever (He et al., 2024) leverages GNNs to enhance LLM reasoning, while GNN-RAG (Mavromatis and Karypis, 2024) employs GNNs for retrieval and LLMs for generation. (4) LLMs + KG: methods that employ LLMs for both retrieval and generation over knowledge graphs. We compare with ToG (Sun et al., 2023), RoG (Luo et al., 2023), KD-CoT (Wang et al., 2023b), and SubGraphRAG (Li et al., 2024).
@@
To ensure a fair comparison, we train each retriever and generator on a single dataset, rather than jointly training across multiple datasets as done in RoG and GNN-RAG. Moreover, to eliminate the influence of backbone model, we use the same Llama-3.1-8B-Instruct model (Touvron et al., 2023) for most LLM-based baselines, including RoG, KD-CoT, SubGraphRAG, G-Retriever, and GNN-RAG. We also evaluate with different backbone models in Appendix D.
@@ 5.2. Overall Performance Comparison
  • The LLM-only models, which lack access to graph information, generally perform worse than methods that incorporate graphs. However, stronger models, such as GPT-4.1, achieve better results, suggesting that large LLMs already possess internal knowledge relevant to the benchmarks. This observation justifies our choice that using the same LLM backbone for baselines to ensure a fair comparison.
@@ 5.4. Ablation Studies
Figure 4. Performance of RoE and its variants pretrained on the WebQSP dataset. Marker shapes denote different model variants, while filled and hollow markers represent Hit and F1 metrics, respectively.
@@ 6. Conclusion
Acknowledgment
This work is supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IIS2504089, DUE2234015, CNS2246050, DRL2405483, IIS 2524379, and NAIRR 250188, the Michigan Department of Agriculture and Rural Development, US Dept of Commerce, Gates Foundation, Amazon Faculty Award, Meta, NVIDIA, Microsoft and SNAP.
@@ Appendix C Experimental Settings
Appendix D Results with different backbone models
To evaluate robustness across model sizes, we additionally train RoE-3B and other baseline models with Llama-3.2-3B-Instruct backbone. All models are trained on WebQSP and evaluated on both WebQSP and MetaQA.
@@ Appendix D Results with different backbone models
Table 4. Performance comparison on the WebQSP and MetaQA datasets. All models use the Llama-3.2-3B-Instruct backbone and are trained on the WebQSP dataset.
WebQSP $\rightarrow$ WebQSP WebQSP $\rightarrow$ MetaQA
Method Hit F1 Hit F1
G-Retrieval-3B 67.32 47.25 12.7 5.58
RoG-3B 73.95 55.79 10.80 6.00
GNN-RAG-3B 81.51 67.16 41.70 13.94
RoE-3B 85.56 70.62 74.20 57.12
From the results in Table 4, RoE continues to outperform all baselines by a large margin, especially in the transfer setting on MetaQA. Notably, RoE-3B surpasses baselines even when they use Llama-8B, especially on the transfer setting. GRPO vs. GSPO yields similar performance, showing that RoE is not sensitive to the RL algorithm choice.